∫e cos^2x .sin2x dx

∫ecos^2x.sin2x dx

Grade:12

1 Answers

Jitender Singh IIT Delhi
9 years ago
Ans:
Hello Student,
Please find answer to your question below

I = \int e^{cos^{2}x}sin2x.dx
cos^{2}x = t
2cosx(-sinx)dx = dt
-sin2xdx = dt
I = \int e^{t}dt
I = e^{t}+c
I = e^{cos^{2}x}+c